# A large Deviation Principle for the Normalized Excursion of $\alpha$-Stable Lévy Processes without Negative Jump

### v2

@@ Abstract
[…]
A large deviation principle for the normalized excursion
of $\alpha$-stable Lévy processes without negative jumps
by
Léo Dort , Christina Goldschmidt & Grégory Miermont
Figure 1: Simulation of the area under the normalized excursion of a $\frac{4}{3}$-stable Lévy process without negative jumps.
@@ 1.   Introduction
Our main goal in this paper is to establish a strong large deviation principle for the normalized bridges and excursions of the process $L$. This contributes to the rich study of functional large deviation principles (LDP) for Lévy processes and related processes, which has its roots in the classical theorem of Cramér and its extension to random walks, see Chapter 5.1 in [9], and references therein. For a Lévy process $(X(t),t\geq 0)$, the natural setting is to consider the family of renormalized processes
@@
The case where $X$ is Brownian motion is addressed by a famous theorem of Schilder [28], who showed a large deviation principle with speed $T$ in the space of continuous functions, where the rate function is the Dirichlet energy. Other Lévy processes are addressed in the landmark paper by Lynch and Sethuraman [20], which was extended in various directions in particular by Borovkov, Mogul’skiĭ and others, see for instance [21, 6, 7, 22, 16, 17] and references therein. However, the vast majority of the results in the above references assume that the Lévy process has a vanishing Gaussian part, as well as the Cramér condition that $\mathbb{E}\left[e^{\lambda X(1)}\right]<\infty$ for every $\lambda$ in a non-empty neighborhood of 00. These conditions imply that the trajectories of the Lévy process have finite variation almost surely, and that the law of $X(1)$ has exponential tails. Various situations may occur when the Cramér condition does not hold. The references [12, 13] consider the case of stretched exponential tails, while [26] considers the situation where the tail of $X(1)$ is regularly varying. In these cases, large deviation principles hold with sublinear speeds, and even logarithmic speed when the tails are regularly varying.
@@
However, considering bridges and excursions of such processes is a way to root them at a given value at time $1$, which prevents the process from “going up too much”. This phenomenon is well-known and was already exploited in [1, 18], in particular in the study of the heights of random trees. However, to our knowledge it has not been used to derive an actual LDP result for stable excursions and bridges, and we fill this gap in the present work. Note that LDPs for bridge-like random walks and Lévy processes were considered under the Cramér condition in [5].
@@
Although we believe our results should have extensions to a much larger class of Lévy process bridges and excursions, we focus here on the particular case of stable processes. In this case, precise estimates are known for the transition densities and the entrance law of the excursion measure, which allow us to provide a rather straightforward extension of the method of proof presented by Serlet in [29]. However, our proof differs from the latter in a crucial aspect. As is often the case in LDP theory, the choice of an appropriate topology on the path space is an important matter. In [20], the authors derived a strong LDP for Lévy processes as in (1.1) in a “weak” topology, and observed that the rate function is not good (in the sense that it does not have compact level sets) in the natural Skorokhod J1 topology. In a series of papers, Borovkov and Mogul’skiĭ improved these results by considering local versions of the LDP in the J1 topology, or by working in the completion of the Skorokhod J1 metric. Unsurprisingly, similar questions arise in our context. Indeed, since it is much less costly for the process $L$ to go up rather than to go down, a similar property also holds for its bridges and excursions, and this implies that in the large deviation regimes considered in this paper (and in stark contrast to [26]), we cannot distinguish between the situation where the process performs one big jump, or two jumps of half the size at extremely close locations, precluding exponential tightness in the J1 topology.
@@
Fortunately, Skorokhod introduced three other possible topologies, called M1, J2 and M2, on spaces of càdlàg functions, and M1 will turn out to suit our purposes, with a very minor adaptation called M1’ that was already considered in [2] and in [30], in the contexts of LDPs for Lévy processes and random walks with Weibull increments, and of a contraction principle for random walk LDPs. We note that Mogul’skiĭ and others [7, 22] also considered similar topologies in the context of processes satisfying the Cramér condition. The M1 topology was also used by O’Brien [24] to prove LDPs for the processes $(L\vee 1)^{\varepsilon}$ as $\varepsilon\downarrow 0$, but these large deviations regimes are very different from the one considered here. We will recall how to define a distance function $\mathrm{dist}$ that induces the M1’ topology and makes $(\mathbb{D}[0,1],\mathrm{dist})$ a Polish space. It is in this space that a strong LDP holds for the excursion and bridge of the process $L$.
@@ 1.1.  Main results
Let $\mathbbm{e}=(\mathbbm{e}_{t},0\leq t\leq 1)$ be the normalized excursion of $L$ above its past infimum [8], and let $\mathbbm{b}^{(x)}=(\mathbbm{b}^{(x)}_{t},0\leq t\leq t)$ be the bridge of $L$ from 00 to $x\in\mathbb{R}$ with unit duration. We set $\mathbbm{b}=\mathbbm{b}^{(0)}$. These processes satisfy a.s. $\mathbbm{e}_{0}=\mathbbm{e}_{1}=\mathbbm{b}^{(x)}_{0}=0$, $\mathbbm{b}^{(x)}_{1}=x$, and $\mathbbm{e}_{t}>0$ for every $t\in(0,1)$. We view $\mathbbm{e}$ and $\mathbbm{b}^{(x)}$ as random variables in the space $\mathbb{D}[0,1]$ of “càdlàg” functions $f:[0,1]\to\mathbb{R}$, that is, functions which are right-continuous at every point $t\in[0,1)$ and have left-limits at every point $t\in(0,1]$. We denote by $f(t+)=f(t)$ and $f(t-)$ the right- and left-limits of $f\in\mathbb{D}[0,1]$ at $t$, whenever applicable. We turn $\mathbb{D}[0,1]$ into a measurable space by equipping it with the $\sigma$-algebra generated by the evaluation maps $f\mapsto f(t)$ for $t\in[0,1]$.
@@
where $(f_{\mathrm{ac}}^{\prime}(s))_{-}$ denotes the negative part of $f_{\mathrm{ac}}^{\prime}(s)$, and we let $I_{\mathbbm{e}}(f)=\infty$ if $f$ is not nonnegative, or if $f$ does not satisfy $f(1)=0$, or if $f$ does not have bounded variation. Here, the term $\infty\cdot f_{\mathrm{sing}\downarrow}(1)$ accounts for the fact that (1.5) yields +$\infty$+ ∞ when $f_{\downarrow}$ is not absolutely continuous. In this way, we note that the shape of the rate function is exactly that involved in [20, Theorem 5.1] (although this theorem does not apply in our context) and the other references mentioned earlier in this introduction.
We defer a discussion of the topology until Section 2.2, where we will introduce a distance $\mathrm{dist}$ on the set $\mathbb{D}[0,1]$. We can now state our main result.
@@
Recall that in a Polish space, an LDP with a good rate function implies exponential tightness (see [20, Lemma 2.6] or [10]). However, since the rate function $I_{\mathbbm{e}}$ is not good, this does not rule out the possibility that $\varepsilon\mathbbm{e}$ satisfies the LDP in the J1 topology, and we do not know whether this property holds or not. Note that this problem has been the topic of several references dealing with random walks and Lévy processes under the Cramér condition, including [20, 21, 6], and at present there is no complete answer to this question in that context either. However, we believe that the M1 topology is a more natural choice in this context, since it is arguably a strong topology for which the rate functions are better behaved.
Theorem 1.2 allows us to deduce general LDPs for functionals of the normalized stable excursion. This extends the results of [11] dealing with Brownian excursions to the case of stable excursions, and was the initial motivation for the present work. Define the sets
@@
Using [14, Theorem 4.5], we have that (1.8) implies the following asymptotics for the Laplace transform and the moments:
@@
Taking as a particular case the functions $\Phi(f)=\int_{0}^{1}f(s)\,\mathrm{d}s$ and $\Phi(f)=\sup_{s\in[0,1]}f(s)$, we obtain the following result, which improves [25, Corollary 1.2] by pinning down the precise constants.
@@ 1.2.  Large deviation principles for bridges
$D_{\mathrm{br}}^{(a)}[0,1]$$\coloneqq\left\{f\in\mathbb{D}[0,1]:\>f(1)=a\right\},$
$H_{\mathrm{br}}^{(a)}$ $\coloneqq H^{(\alpha)}\cap D_{\mathrm{br}}^{(a)}[0,1].$
@@
Let $(a_{\varepsilon})_{\varepsilon>0}$ be such that $\varepsilon a_{\varepsilon}\to a$ as $\varepsilon\to 0$. Then the laws of $(\varepsilon\mathbbm{b}^{(a_{\varepsilon})}_{t})_{t\in[0,1]}$ satisfy a LDP in $(\mathbb{D}[0,1],\mathrm{dist})$ as $\varepsilon\downarrow 0$ with speed $\varepsilon^{-\alpha^{\prime}}$ and good rate function $I_{\mathbbm{b},a}$ defined by
@@
$K_{\mathrm{br}}\coloneqq K^{(\alpha)}\cap D_{\mathrm{br}}^{(0)}[0,1].$
@@
$\gamma_{\Phi}\coloneqq\max\left\{\Phi(f):\>f\in K_{\mathrm{br}}\right\}.$
@@
As an application, using the same proof as for 1.6, we may reprove an exact logarithmic asymptotic for the right tails of the supremum of the stable Lévy bridge obtained by Kortchemski in [18, Corollary 13].
Corollary 1.9 ([18, Corollary 13]).
@@ 1.3.  Outline of the proofs and organization of the paper
The proofs for excursions and bridges are very much alike, but some extra technicalities arise for excursions, so we focus mostly on this case, and deal with bridges in Section 7. Section 2 will recall the basics of stable processes, bridges and excursions, as well as the results on the M1 topology that will be needed in this paper.
In order to prove 1.2, we first establish the LDP for the finite-dimensional marginals of $\varepsilon\mathbbm{e}$ (3.1). This relies on the explicit form of the finite-dimensional marginals of the stable excursion in terms of stable densities and related quantities, which is recalled in Section 2.1. The key input is the following estimate for stable densities $p_{t}(x)=\mathbb{P}\left(L_{t}\in\,\mathrm{d}x\right)/\,\mathrm{d}x$, [27, Equation (14.35)]),
@@
The asymmetry of these two asymptotic behaviors will play a key role. This will imply that $\varepsilon\mathbbm{e}$ satisfies a large deviation principle for the weak topology on $\mathbb{D}[0,1]$ of pointwise convergence at continuity points of the limit. This result is proved in Section 5, which is also devoted to the identification of the rate function. In order to prove an LDP in $(\mathbb{D}[0,1],\mathrm{dist})$, we show that the laws of $(\varepsilon\mathbbm{e},\varepsilon\in(0,1))$ are exponentially tight in this space. This relies on a Kolmogorov-type criterion which we prove in Section 2.2.3, and apply to our present context is Section 4. Finally, 1.4 is proved in Section 6.1 using the ideas of Fill and Janson [11], who treated the Brownian case.
Acknowledgements
Thanks are due to Loïc Chaumont for an interesting conversation around stable excursions. We also thank the anonymous referee for their careful reading of the paper, and for mentioning a number of relevant references. CG’s research was supported by EPSRC Fellowship EP/N004833/1.
@@ 2.1.  Excursions and bridges of stable Lévy processes without negative jumps
We begin by recalling the definitions of stable excursions and bridges, for which we mainly refer to [8]. We denote by $\mathbb{P}_{x}$ the law under which the canonical càdlàg process $(L_{t},t\geq 0)$ is a stable Lévy process without negative jumps with exponent $\alpha$, started from $x$, and we set $\mathbb{P}=\mathbb{P}_{0}$. We let $(\mathcal{F}_{t})_{t\geq 0}$ be the natural filtration. We denote by $(p_{t})_{t\geq 0}$ the continuous transition semigroup density of $L$ under $\mathbb{P}_{x}$, which possesses the scaling property
@@
Denote by $\underline{n}$ the Itô measure of $L$ above its past infimum. The $\sigma$-finite “law” of the lifetime under $\underline{n}$ has been calculated by Monrad and Silverstein (see [23, Lemma 3.2])
@@
$\mathbb{P}_{x}^{(0,\infty)}\left(A,\,t<\zeta\right)\coloneqq\mathbb{P}\left(A,\,t<\tau_{(-\infty,0)}\right),\quad t\geq 0,\quad A\in\mathcal{F}_{t}.$
@@
We denote by $(q_{x}(t))_{t\geq 0}$ the density of the first hitting time of 00 under $\mathbb{P}_{x}^{(0,\infty)}$. Thanks to the absence of negative jumps, the density $(q_{x}(t))_{t\geq 0}$ can be related to the law of $L$ through the relation (see for instance [3, Corollary VII.3])
@@
(see [23, Lemma 3.2]). Combining the above expressions, we may show that the law of the normalized excursion has a density with respect to the Lebesgue measure. Indeed, if $f:\mathbb{R}\rightarrow\mathbb{R}_{+}^{*}$ and $g:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ are two nonnegative measurable functions, we then have

### v3

# A large deviation principle for the normalized excursion of an $\alpha$-stable Lévy process without negative jumps

@@ Abstract.
Key words and phrases:
Large deviations principles, $\alpha$-stable Lévy processes, excursions and bridges, Skorokhod topologies
2010 Mathematics Subject Classification:
60G52, 60F10, 60G51, 60G17
The second author was supported by EPSRC Fellowship EP/N004833/1.
Simulation of the area under the normalized excursion of a $\frac{4}{3}$-stable Lévy process without negative jumps.
@@ 1. Introduction
Our main goal in this paper is to establish a strong large deviation principle for the normalized bridges and excursions of the process $L$. This contributes to the rich study of functional large deviation principles (LDP) for Lévy processes and related processes, which has its roots in the classical theorem of Cramér and its extension to random walks, see Chapter 5.1 in Dembo and Zeitouni (1998), and references therein. For a Lévy process $(X(t),t\geq 0)$, the natural setting is to consider the family of renormalized processes
@@
The case where $X$ is Brownian motion is addressed by a famous theorem of Schilder (1966), who showed a large deviation principle with speed $T$ in the space of continuous functions, where the rate function is the Dirichlet energy. Other Lévy processes are addressed in the landmark paper by Lynch and Sethuraman (1987), which was extended in various directions in particular by Borovkov, Mogul’skiĭ and others (see, for instance, Mogul’skiĭ (1993); Borovkov and Mogul’skiĭ (2013b, 2014); Mogul’skiĭ (2017); Klebaner and Mogul’skiĭ (2019); Klebaner et al. (2020) and references therein). However, the vast majority of the results in the above references assume that the Lévy process has a vanishing Gaussian part, as well as the Cramér condition that $\mathbb{E}\left[e^{\lambda X(1)}\right]<\infty$ for every $\lambda$ in a non-empty neighborhood of 00. These conditions imply that the trajectories of the Lévy process have finite variation almost surely, and that the law of $X(1)$ has exponential tails. Various situations may occur when the Cramér condition does not hold. The references Gantert (1998); Gantert et al. (2014) consider the case of stretched exponential tails, while Rhee et al. (2019) considers the situation where the tail of $X(1)$ is regularly varying. In these cases, large deviation principles hold with sublinear speeds, and even logarithmic speed when the tails are regularly varying.
@@
However, considering bridges and excursions of such processes is a way to root them at a given value at time $1$, which prevents the process from “going up too much”. This phenomenon is well-known and was already exploited in Addario-Berry et al. (2013); Kortchemski (2017), in particular in the study of the heights of random trees. However, to our knowledge it has not been used to derive an actual LDP result for stable excursions and bridges, and we fill this gap in the present work. Note that LDPs for bridge-like random walks and Lévy processes were considered under the Cramér condition in Borovkov and Mogul’skiĭ (2013a).
@@
Although we believe our results should have extensions to a much larger class of Lévy process bridges and excursions, we focus here on the particular case of stable processes. In this case, precise estimates are known for the transition densities and the entrance law of the excursion measure, which allow us to provide a rather straightforward extension of the method of proof presented by Serlet (1997). However, our proof differs from the latter in a crucial aspect. As is often the case in LDP theory, the choice of an appropriate topology on the path space is an important matter. In Lynch and Sethuraman (1987), the authors derived a strong LDP for Lévy processes as in (1.1) in a “weak” topology, and observed that the rate function is not good (in the sense that it does not have compact level sets) in the natural Skorokhod J1 topology. In a series of papers, Borovkov and Mogul’skiĭ improved these results by considering local versions of the LDP in the J1 topology, or by working in the completion of the Skorokhod J1 metric. Unsurprisingly, similar questions arise in our context. Indeed, since it is much less costly for the process $L$ to go up rather than to go down, a similar property also holds for its bridges and excursions, and this implies that in the large deviation regimes considered in this paper (and in stark contrast to Rhee et al. (2019)), we cannot distinguish between the situation where the process performs one big jump, or two jumps of half the size at extremely close locations, precluding exponential tightness in the J1 topology.
@@
Fortunately, Skorokhod introduced three other possible topologies, called M1, J2 and M2, on spaces of càdlàg functions, and M1 will turn out to suit our purposes, with a very minor adaptation called M1’ that was already considered in Bazhba et al. (2020) and in Vysotsky (2021), in the contexts of LDPs for Lévy processes and random walks with Weibull increments, and of a contraction principle for random walk LDPs. We note that Mogul’skiĭ and others (Borovkov and Mogul’skiĭ (2014); Mogul’skiĭ (2017)) also considered similar topologies in the context of processes satisfying the Cramér condition. The M1 topology was also used by O’Brien (1999) to prove LDPs for the processes $(L\vee 1)^{\varepsilon}$ as $\varepsilon\downarrow 0$, but these large deviations regimes are very different from the one considered here. We will recall how to define a distance function $\mathrm{dist}$ that induces the M1’ topology and makes $(\mathbb{D}[0,1],\mathrm{dist})$ a Polish space. It is in this space that a strong LDP holds for the excursion and bridge of the process $L$.
@@ 1.1. Main results
Let $\mathbbm{e}=(\mathbbm{e}_{t},0\leq t\leq 1)$ be the normalized excursion of $L$ above its past infimum (see Chaumont (1997)), and let $\mathbbm{b}^{(x)}=(\mathbbm{b}^{(x)}_{t},0\leq t\leq t)$ be the bridge of $L$ from 00 to $x\in\mathbb{R}$ with unit duration. We set $\mathbbm{b}=\mathbbm{b}^{(0)}$. These processes satisfy a.s. $\mathbbm{e}_{0}=\mathbbm{e}_{1}=\mathbbm{b}^{(x)}_{0}=0$, $\mathbbm{b}^{(x)}_{1}=x$, and $\mathbbm{e}_{t}>0$ for every $t\in(0,1)$. We view $\mathbbm{e}$ and $\mathbbm{b}^{(x)}$ as random variables in the space $\mathbb{D}[0,1]$ of “càdlàg” functions $f:[0,1]\to\mathbb{R}$, that is, functions which are right-continuous at every point $t\in[0,1)$ and have left-limits at every point $t\in(0,1]$. We denote by $f(t+)=f(t)$ and $f(t-)$ the right- and left-limits of $f\in\mathbb{D}[0,1]$ at $t$, whenever applicable. We turn $\mathbb{D}[0,1]$ into a measurable space by equipping it with the $\sigma$-algebra generated by the evaluation maps $f\mapsto f(t)$ for $t\in[0,1]$.
@@
where $(f_{\mathrm{ac}}^{\prime}(s))_{-}$ denotes the negative part of $f_{\mathrm{ac}}^{\prime}(s)$, and we let $I_{\mathbbm{e}}(f)=\infty$ if $f$ is not nonnegative, or if $f$ does not satisfy $f(1)=0$, or if $f$ does not have bounded variation. Here, the term $\infty\cdot f_{\mathrm{sing}\downarrow}(1)$ accounts for the fact that (1.5) yields +$\infty$+ ∞ when $f_{\downarrow}$ is not absolutely continuous. In this way, we note that the shape of the rate function is exactly that involved in (Lynch and Sethuraman, 1987, Theorem 5.1) (although this theorem does not apply in our context) and the other references mentioned earlier in this introduction.
We defer a discussion of the topology until Section 2.2, where we will introduce the distance $\mathrm{dist}$ on the set $\mathbb{D}[0,1]$. We can now state our main result.
@@
Recall that in a Polish space, an LDP with a good rate function implies exponential tightness (see (Lynch and Sethuraman, 1987, Lemma 2.6) or Eichelsbacher and Grunwald (1999)). However, since the rate function $I_{\mathbbm{e}}$ is not good, this does not rule out the possibility that $\varepsilon\mathbbm{e}$ satisfies the LDP in the J1 topology, and we do not know whether this property holds or not. Note that this problem has been the topic of several references dealing with random walks and Lévy processes under the Cramér condition, including Lynch and Sethuraman (1987); Mogul’skiĭ (1993); Borovkov and Mogul’skiĭ (2013b), and at present there is no complete answer to this question in that context either. However, we believe that the M1 topology is a more natural choice in this context, since it is arguably a strong topology for which the rate functions are better behaved.
Theorem 1.2 allows us to deduce general LDPs for functionals of the normalized stable excursion. This extends the results of Fill and Janson (2009) dealing with Brownian excursions to the case of stable excursions, and was the initial motivation for the present work. Define the sets
@@
Using (Janson and Chassaing, 2004, Theorem 4.5), we have that (1.8) implies the following asymptotics for the Laplace transform and the moments:
@@
Taking as a particular case the functions $\Phi(f)=\int_{0}^{1}f(s)\,\mathrm{d}s$ and $\Phi(f)=\sup_{s\in[0,1]}f(s)$, we obtain the following result, which improves (Profeta, 2021, Corollary 1.2) by pinning down the precise constants.
@@ 1.2. Large deviation principles for bridges
$D_{\mathrm{br}}^{(a)}[0,1]$ =$=\left\{f\in\mathbb{D}[0,1]:\>f(1)=a\right\},$
$H_{\mathrm{br}}^{(a)}$ $=H^{(\alpha)}\cap D_{\mathrm{br}}^{(a)}[0,1].$
@@
Let $(a_{\varepsilon})_{\varepsilon>0}$ be such that $\varepsilon a_{\varepsilon}\to a$ as $\varepsilon\to 0$. Then the laws of $(\varepsilon\mathbbm{b}^{(a_{\varepsilon})}_{t})_{t\in[0,1]}$ satisfy an LDP in $(\mathbb{D}[0,1],\mathrm{dist})$ as $\varepsilon\downarrow 0$ with speed $\varepsilon^{-\alpha^{\prime}}$ and good rate function $I_{\mathbbm{b},a}$ defined by
@@
$K_{\mathrm{br}}=K^{(\alpha)}\cap D_{\mathrm{br}}^{(0)}[0,1].$
@@
$\gamma_{\Phi}=\max\left\{\Phi(f):\>f\in K_{\mathrm{br}}\right\}.$
@@
As an application, using the same proof as for Corollary 1.6, we may reprove an exact logarithmic asymptotic for the right tails of the supremum of the stable Lévy bridge obtained by Kortchemski (2017).
Corollary 1.9 ((Kortchemski, 2017, Corollary 13)).
@@ 1.3. Outline of the proofs and organization of the paper
The proofs for excursions and bridges are very much alike, but some extra technicalities arise for excursions, so we focus mostly on that case, and deal with bridges in Section 7. Section 2 will recall the basics of stable processes, bridges and excursions, as well as the results on the M1 topology that will be needed in this paper.
In order to prove Theorem 1.2, we first establish the LDP for the finite-dimensional marginals of $\varepsilon\mathbbm{e}$ (Proposition 3.1). This relies on the explicit form of the finite-dimensional marginals of the stable excursion in terms of stable densities and related quantities, which is recalled in Section 2.1. The key input is the following estimate for stable densities $p_{t}(x)=\mathbb{P}\left(L_{t}\in\,\mathrm{d}x\right)/\,\mathrm{d}x$ (see (Sato, 1999, Equation (14.35))),
@@
The asymmetry of these two asymptotic behaviors will play a key role. This will imply that $\varepsilon\mathbbm{e}$ satisfies a large deviation principle for the weak topology on $\mathbb{D}[0,1]$ of pointwise convergence at continuity points of the limit. This result is proved in Section 5, which is also devoted to the identification of the rate function. In order to prove an LDP in $(\mathbb{D}[0,1],\mathrm{dist})$, we show that the laws of $(\varepsilon\mathbbm{e},\varepsilon\in(0,1))$ are exponentially tight in this space. This relies on a Kolmogorov-type criterion which we prove in Section 2.2.3, and apply to our present context in Section 4. Finally, Theorem 1.4 is proved in Section 6.1 using the ideas of Fill and Janson (2009), who treated the Brownian case.
@@ 2.1. Excursions and bridges of stable Lévy processes without negative jumps
We begin by recalling the definitions of stable excursions and bridges, for which we mainly refer to Chaumont (1997). We denote by $\mathbb{P}_{x}$ the law under which the canonical càdlàg process $(L_{t},t\geq 0)$ is a stable Lévy process without negative jumps with exponent $\alpha$, started from $x$, and we set $\mathbb{P}=\mathbb{P}_{0}$. We let $(\mathcal{F}_{t})_{t\geq 0}$ be the natural filtration. We denote by $(p_{t})_{t\geq 0}$ the continuous transition semigroup density of $L$ under $\mathbb{P}_{x}$, which possesses the scaling property
@@
Denote by $\underline{n}$ the Itô measure of $L$ above its past infimum. The $\sigma$-finite “law” of the lifetime under $\underline{n}$ has been calculated by (Monrad and Silverstein, 1979, Lemma 3.1):
@@
$\mathbb{P}_{x}^{(0,\infty)}\left(A,\,t<\zeta\right)=\mathbb{P}\left(A,\,t<\tau_{(-\infty,0)}\right),\quad t\geq 0,\quad A\in\mathcal{F}_{t}.$
@@
We denote by $(q_{x}(t))_{t\geq 0}$ the density of the first hitting time of 00 under $\mathbb{P}_{x}^{(0,\infty)}$. Thanks to the absence of negative jumps, the density $(q_{x}(t))_{t\geq 0}$ can be related to the law of $L$ as follows (see, for instance, (Bertoin, 1996, Corollary VII.3)):
@@
(see (Monrad and Silverstein, 1979, Lemma 3.2)). Combining the above expressions, we may show that the law of the normalized excursion has a density with respect to the Lebesgue measure. Indeed, if $f:\mathbb{R}\rightarrow\mathbb{R}_{+}^{*}$ and $g:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ are two nonnegative measurable functions, we then have